3 square root of 2 (8 plus square root 5) Question 2: 7 (negative 2 square root of 5 plus 6) Question 3: 2 square root of X minus 3 square r

Algebra ->  Square-cubic-other-roots -> SOLUTION: Simplify: Question 1: 3 square root of 2 (8 plus square root 5) Question 2: 7 (negative 2 square root of 5 plus 6) Question 3: 2 square root of X minus 3 square r      Log On


   



Question 422265: Simplify:
Question 1:
3 square root of 2 (8 plus square root 5)
Question 2:
7 (negative 2 square root of 5 plus 6)
Question 3:
2 square root of X minus 3 square root of x plus 4 square root of x

Answer by shree840(260)

I will put values sq root2=1.41
1.
3*1.41(8+2.23) sq rt 5=2.23
43.27
2. 7*(-2sq root 5 +6)
= 7*(-2*2.23+6)
=7*(1.54)=10.78
3. 2sqrootx-3sqrootx+4sqrootx
add or subtract
(2+4-3)sqroots
3sqrootx